A 10kg block is placed on 40 kg block.a force of 200N is applied on small block.If  coefficient of static and kinetic friction are 0.4 and 0.6 respectively.then find the accelaration of the system.

Grade:12

2 Answers

sonika p
13 years ago

where is the friction acting between the blocks or between block and ground. check ur question how can coefficient of kinetic friction greater than static

vikas askiitian expert
13 years ago

at upper block  ,static friction force is less than force applied

  so the kinetic friction now acts on the block

 at upper block  ,

                         F-uk10g =10a

 accleration of small block           a= 200-60/10=14

  now this friction force becomes driving force for the bigger block

     accleration of bigger block  a1=60/50=1.2